Ask
your own question, for FREE!
Mathematics
9 Online
MEDAL AND FAN Hitomi has a bag with four chips in it numbered 1 to 4. She randomly draws a chip and replaces it three times. What is the probability that Hitomi will draw a 1 all three times?
Still Need Help?
Join the QuestionCove community and study together with friends!
@jim_thompson5910
what is the probability of drawing a 1?
1/infinity
:)
Idk please help lol @jim_thompson5910
Still Need Help?
Join the QuestionCove community and study together with friends!
how many tiles have "1" on it? how many tiles are there total?
only 1 and there are 4
so the probability of drawing a "1" is 1/4 = 0.25
the probability of drawing "1" three times (with replacement) in a row is (1/4)^3 = ???
0.015625
Still Need Help?
Join the QuestionCove community and study together with friends!
yes that's the probability in decimal form
how about in fraction form?
Idk
|dw:1432082988748:dw|
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
17 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
3 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal